A vessel contains oil (density =0.8gm/cm3) over mercury (density=13.6 gm cm3 ). A homogeneous sphere floats with half its volume immersed in mercury and the other half in oil. The density of the material of the sphere in gm/cm3
(a) 3.3
(b) 6.4
(c) 7.2
(d) 12.8

Grade:12th pass

1 Answers

Aditi Chauhan
9 years ago

(c) Weight of sphere

= Upthrust due to Hg + Up + Upthrust due to oil

Vdg = V/2dHgg + V/doil * g

? d = dHg + doil/2 = 13.6 + 0.8/2 = 7.2g/cm3